OpenStudy (anonymous):

Two of your friends, Matt and Karen, both run to you to settle a dispute. They were working on a math problem, and got different answers. Wisely, you decide to look at their work to see if you can spot the source of confusion. Matt Karen 6 – 4(3 – 5)2 + 30 ÷ 5 6 – 4(3 – 5)2 + 30 ÷ 5 6 – 4(–2)2 + 30 ÷ 5 6 – 4(–2)2 + 30 ÷ 5 6 – 4(4) + 30 ÷ 5

OpenStudy (anonymous):

hey nansie are you there?

OpenStudy (anonymous):

yeah, im here.

OpenStudy (anonymous):

ok so are you just trying to find what went wrong in both persons answers? it seems like the question isn't quite finished

OpenStudy (anonymous):

yeah. the question got cut off. ill post up they're problems

OpenStudy (anonymous):

ok gimme a moment here

OpenStudy (anonymous):

that wasnt much help oops

OpenStudy (anonymous):

whoa wait they were just here!

OpenStudy (anonymous):



OpenStudy (anonymous):

let me know if that worked?

OpenStudy (anonymous):

it worked just lemme work on the problem

OpenStudy (anonymous):

wait its sorta loading... whatever you did before this worked though

OpenStudy (anonymous):

Matt Karen 6 – 4(3 – 5)2 + 30 ÷ 5 6 – 4(3 – 5)2 + 30 ÷ 5 6 – 4(–2)2 + 30 ÷ 5 6 – 4(–2)2 + 30 ÷ 5 6 – 4(4) + 30 ÷ 5 6 – 4(−4) + 30 ÷ 5 6 – 16 + 30 ÷ 5 6 + 16 + 30 ÷ 5 −10 + 30 ÷ 5 6 + 16 + 6 20 ÷ 5 22 + 6 4 28

OpenStudy (anonymous):

you dont see all the question marks ??

OpenStudy (anonymous):

ok now i got it! gimme a sec here..

OpenStudy (anonymous):

matt's answer is wrong because in the third line its supposed to be -4 because 2*-2 does not equal 4

OpenStudy (anonymous):

so i was right??, cause i explained how he didnt use pemdas .

OpenStudy (anonymous):

yep matt was right not because he didn't use pemdas but the fact that he calculated wrong

OpenStudy (anonymous):

can you help me with another one, that i didnt understand too well?

OpenStudy (anonymous):

no problem fire away

OpenStudy (anonymous):

Ellen works for a high-speed rail company that wants to develop a new rail line. Ellen’s project is to find a train that is the second fastest in the world. The Shinkansen Bullet Train in Japan is reported to go as fast as 320 kilometers per hour. The TGV train in France can reach speeds of 89.44 meters per second. Explain to Ellen how to find what speed her new train must go to be the second fastest in this group. Then, find an appropriate speed for Ellen’s train in miles per hour and use the formula distance = speed • time to find how far the train can go in 2.5 hours. Use complete sentences and show your work.

OpenStudy (anonymous):

Convert all speed to mi / hr Japanese train is 320 km / hr 1mi = 1.6093km or 1 km = 1/1.6093 mi So 320 km / hr = (320) (1/1.6093) mi / hr = 198.8 mi / hr French train is 89.44 meters per second. Convert it first to meter per hour by multiplying by 3600 as there are 3600 seconds in an hour 89.44 x 3600 = 321984 meters per hour Convert meters to km by dividing by 1000 as there are 1000 meters in a kilometer 321984 / 1000 = 321.984 km / hr convert to miles per hour by dividing by 1.6093 321.984 / 1.6093 = 200.1 mi / hr So the fastest train is the French train at 200.1 mi / hr The Japanese train is at 198.8 mi / hr To be the second fastest train, the speed has to be between those two numbers. I will take a midpoint (or average) between those two numbers: Ellen's train must be at speed (200.1 + 198.8)/2 = 199.5 mi / hr to be the second fastest train. In 2.5 hours the train will travel: 199.5 x 2.5 = 499 miles.

OpenStudy (anonymous):

wait did you just answer it for me ?

OpenStudy (anonymous):

was i not supposed to ? oops sorry

OpenStudy (anonymous):

no i was all confused but thanks a lot. i'd love to ask for help again but i guess i shouldn't xD

OpenStudy (anonymous):

no no no plz ask away!

OpenStudy (anonymous):

alright alright

OpenStudy (anonymous):

Beatriz is creating graphs for her city’s Parks Department. Beatriz is trying to persuade the Parks Department to put in more gazebos for shade. For a local park, Beatriz wants to demonstrate the attendance as it changes every week. She also wants to show the temperature changes per month. Sometimes temperatures reach below zero. Explain to Beatriz how to make the graphs correctly based on the data that she will collect and what labels, scales, and intervals each axis must have on each graph. Use complete sentences and support your reasoning. Feel free to generate sample data to help support your response.

OpenStudy (anonymous):

changes every week. changes per month You will notice that week and month are both subdivisions of a year ......................... Graph 1 (24 by 24) Title: Weekly attendance at local park X-axis label: Weeks X-axis scale: Weeks X-axis intervals: 5 weeks Y-axis label: People Y-axis scale: Number of people Y-axis intervals: 300 persons Graph 2 (24 by 24) Title: Temperature throughout two years at local park X-axis label: Months (Starting at "Z" month) X-axis scale: Months X-axis intervals: 1 month Y-axis label: Temperature Y-axis scale: Fahrenheit Y-axis intervals: 10 degrees (maximum: 120 degrees) The two graphs are similar in size so that the presenter can somwhat persuade a relation between temperature, time, and people in the park. Fahrenheit is preferred because the data points will not scatter too far from each other, which would probably occur using Celsius; and will make a linear relation somewhat apparent or more believable. Although temperatures drop below zero, the neccesity of a study for at least two years would ensure that some trend of temperature rise exists between similar months of advancing years. Even though December temperatures are below freezing, the temperature of the previous year's December was more colder than the temperature of that month in this year. Moreover, the presentor can try to relate the data to the common concept of global warming. The number of people attending parks is usually dismal in cities; but events attract crowds, so over 100 humans could be expected in a day or 300 heads can be expected in a week, assuming that a maximum of 3 major events are restricted to take place on Fridays, Saturdays, and either on Sundays or one other weekday. Note that the number of people attending per week will be displayed in subintervals.

OpenStudy (anonymous):

so uhm, should i just copy and paste all of that or should i read it

OpenStudy (anonymous):

you should read it first ;)

OpenStudy (anonymous):

whether it was right or wrong i submitted it thanks so much

OpenStudy (anonymous):

np glad i could help!